In 1993, the life expectancy of males in a certain country was 69.1 years. In 2000,it was 71.7 years. Let e represent the life expectancy in year t and let t represent the numbers of years since 1993.
The linear function E(t)that fits the data is
E(t)=
Use the function to predict the life expectancy of males in 2005
E(12)=

Answer by stanbon(75887)

In 1993, the life expectancy of males in a certain country was 69.1 years.
(0,69.1)
-----------------
In 2000,it was 71.7 years.
(7,71.7)
-----------------
slope = (71.7-69.1)/(7-0) = 2.6/7 = 13/35
intercept = E(0) = 69.1
----------
Let E represent the life expectancy in year t and let t represent the numbers of years since 1993.
The linear function E(t)that fits the data is
E(t)= (13/35)t + 69.1
-------------------------------------------
Use the function to predict the life expectancy of males in 2005
E(12)= (13/35)12+69.1
E(12) = 73.56 years
=============================
Cheers,
Stan H.